An ac series circuit has an impedance of 176 ? , and the phase angle between the current and the voltage of the generator is ? = -68
Answer #1

Current lags, so this is an inductive circuit.

Z = sqrt(R
Answer #2

Current lags, so this is an inductive circuit.

Z = ?(R
Answer #3

phase angle is negative so its and inductor .
Inductance = XL
and tan68 = XL / R
XL = Rtan68 = 2.475R
and as we knoe Impedance Z = sqrt( R^2 + XL^2 )
176^2 = R^2 + 2.475R^2
R = 65.93 Ohm ..............Ans (a)
XL = 2.475 x 65.93 = 163.18 Ohm .........ans(b)